OpenStudy (anonymous):

Divide: (7 + 4i) ÷ (2 + 3i)

hartnn (hartnn):

multiply and divide by the conjugate of the denominator you know whats the conjugate of 2+3i ?

OpenStudy (anonymous):

2+3i?

OpenStudy (anonymous):

right?

hartnn (hartnn):

the conjugate of a+bi will be a-bi what about 2+3i ?

OpenStudy (anonymous):

i d k??

hartnn (hartnn):

just flip the sign of imaginary part! conjugate of 2+3i is 2-3i ok ?

OpenStudy (anonymous):

oooohhh i get it you just flipped the sign

hartnn (hartnn):

yes, now multiply and divide that , what u get ?

OpenStudy (anonymous):

i

OpenStudy (anonymous):

?

OpenStudy (anonymous):

imm sorry i dk what i am doing

hartnn (hartnn):

do this: \(\large \dfrac{(7+4i)(2+3i)}{(2+3i)(2-3i)}\) you can multiply, right ? :P

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do i use the foil method

hartnn (hartnn):

sorry, a typo, do this *** \(\large \dfrac{(7+4i)(2-3i)}{(2+3i)(2-3i)}\) yes, exactly, the foil ! in numerator and denominator both

OpenStudy (anonymous):

ok i got -12i^2-13i+14/9i^2+4

hartnn (hartnn):

numerator is absolutely correct! good the denominator should have -9i^2+4, right ??

OpenStudy (anonymous):

no -9i^-12i+4 i forgot to add -6i+-6i

hartnn (hartnn):

-6i+6i should be there, check again ? and that = 0

OpenStudy (anonymous):

ohhh ok i get it

hartnn (hartnn):

cool, now whats i^2 = .... ? you know ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is it an imaginary number

hartnn (hartnn):

this question belongs o imaginary nu,mber topic, right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what do i do after that

hartnn (hartnn):

so, \(i=\sqrt{-1}\) square both sides of this, what u get ?

OpenStudy (anonymous):

√-12i+1?

OpenStudy (anonymous):

on the numerator

hartnn (hartnn):

what ?? just square both sides of \(i=\sqrt{-1}\) so that we get the value of i^2 which we will substitute in the expression you got....